OpenStudy (anonymous):

A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

Anyone please help?

OpenStudy (anonymous):

 So this is basically what we have here, right?

OpenStudy (anonymous):

i think so it does not give me a picture

OpenStudy (anonymous):

tan30= x/41

OpenStudy (anonymous):

okay so Tan stands for Tangent right? sorry im so bad at math

OpenStudy (anonymous):

-6.40533119665 is what i got when i typed in tan30 so thats correct? I dont understand how you got 41

OpenStudy (anonymous):

Right. If you type tan30 in a calc it comes up as about 0.58. so 0.58=x/41, try getting x by itself

OpenStudy (anonymous):

x=23.78?

OpenStudy (anonymous):

Bingo! So that's the answer to question A

OpenStudy (anonymous):

excellent!! okay now to b :)

OpenStudy (anonymous):

According to SohCahToa, what's Cos90 =?

OpenStudy (anonymous):

-0.44807361612

OpenStudy (anonymous):

but the 90 degrees is negative does that make a diffrence?

OpenStudy (anonymous):

Cosine is adjacent over hypotenuse so it would be cos90 = 23.78/y

OpenStudy (anonymous):

Ohrlly? I don't know how to do negative angles D:

OpenStudy (anonymous):

hmmm, i have no idea either but to do it the way you said i would have to get y by itself so i would multiply 23.78 by itself and then do the same to -90?

OpenStudy (anonymous):

lol the negatives throw me off, let's try tagging someone else xD

OpenStudy (anonymous):

okay =)

OpenStudy (anonymous):

oh hello there!

OpenStudy (anonymous):

Hello :)

OpenStudy (anonymous):

;(

OpenStudy (anonymous):